Weekly Notes - 28th October 2021

With calving finishing last week we are straight into preparation for mating which starts on the 1st of November. Decisions this week have continued to focus on managing surplus pasture, achieving residuals and fully feeding cows as they head into mating.

Key Decisions:

• 7 paddocks have been stepped over for conservation (2 in LI FB, 2 in Std FB, 2 in Std Kale and 1 in LI Kale) and these will be mowed for baleage next week.

• Discretionary topping will be used for individual paddocks that don’t hit residual targets; a flow-on effect from wet conditions and poor utilisation in the previous round.

• all supplementary feed from the four farmlets has been removed

• Cows below BCS 4 are being milked OAD and those from the kale herds get priority in-shed feeding. There are currently x10 Std kale, x10 LI kale, x15 Std FB and x5 LI FB in the OAD group at the moment.

• a 25-day round length means that the grazing plan for the next week has a mixture of 2 or 3 feeds per paddock depending on pre-graze residual.

• For ease of management for the paddocks that will only be getting 2 feeds per day (24-hour grazing) no fence will be put up. However, for paddocks requiring 3 feeds, the farm team will be required to put up on break fence so the cows will have 1x 12 and 1x 24-hour grazing.

• Groundwork is well underway with an Italian being sown today for baleage and grass wintering next winter. Ploughing has also started this week for paddocks going back into permanent pastures.

• Baleage has been cut at the support block and there are more paddocks booked in to be mowed next week.

• the decision has been made to sell the fodder beet that is in the bunker (around 1/3 of a bunker). However, we still have a pile in the paddock that we will keep and use as required for the FB herds.

General Notes:

• We put CIDR’s into cows that had not cycled but were more than 50 days post-calving. Cows were identified from data extracted from the Allflex collars and weekly pre-mating heat recordings using tail paint. The team are excited to have the collar data to assist with heat detection going forward. After consultation with the vets, 46 of the 71 cows identified had a CIDR put in.

As shown in the table below the number of cows that ended up with a CIDR was highest in the Std Kales (18) and Std FB (14), with only 7 CIDR’s being required in both the LI Kale and LI FB herds.

• N will continue to be applied (25 kg N/ha) to the Std farmlets following grazing

• Maintenance fertiliser applications are continuing

• All calves are now outside enjoying this great weather. There are only a handful of calves left on the platform which will be moved to the support block in the coming weeks.

• This week, we have had the digger in emptying out the weeping wall and the solids are being spread over the paddocks that will be going into winter crop next year. Manure samples will be taken to determine the nutrient content in the solids, and then fertilizer application rates can be altered accordingly if need be.

• All four herds were drenched with TopLine pour on (25mls) primarily for lice control last week as there has been lots of scratching going on.

Click the link below to download our full weekly notes